OpenStudy (anonymous):

Which mixed number is equivalent to this improper fraction? 23/6 A. 2 1/2 B. 3 1/6 C. 3 1/2 D. 3 5/6

OpenStudy (alexandervonhumboldt2):

first you need to decide how many times 6 will fit in 23. so you divide 23 by 6 and get 3 wholes and a remainder: 3*6=18. to find the remainder subtract 23-18=5 the remainder is 5. thus the mixed number is \[3\frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

Thank you!

OpenStudy (alexandervonhumboldt2):

\[a \frac{ b }{ c }\] a is the whole number, b is the remainder and c is the denominator. your welcome